# Convergence Rates of Constrained Expected Improvement

### v1

@@ 3.1.1 Improved simple regret upper bound under frequentist assumptions
The proof follows similar steps to that of Theorem 3.4 but further bounds $\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$ using $\gamma_{t}^{f}$. To do so, we first recognize that the bound using $\gamma_{t}^{f}$ (Lemma A.3) is established in the noisy case where the posterior standard deviation has a different form as in (19). Using Lemma A.4, we can establish that the noise-free posterior standard deviation also satisfies $\sum_{i=0}^{t-1}\sigma_{i}^{f}({\bm{x}}_{i+1})\leq\sqrt{C_{\gamma}t\gamma_{t}}$. Then, from Lemma A.5, we can find a small enough $\sigma_{i}({\bm{x}}_{i+1})$. Specifically, choose $k=[t/3]$, where $[x]$ denotes the largest integer smaller than $x$. Thus, we have $3k\leq t\leq 3(k+1)$. Then, there exists $k\leq t_{k}\leq 3k$ such that $f^{+}_{t_{k}}-f^{+}_{t_{k}+1}\leq\frac{2B_{f}}{k}$ and $\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})\leq\frac{\sqrt{t\gamma_{t}}}{k}$. The rest of the proof follows from that of Theorem 3.8.
@@
As mentioned above, the rates in Theorem 3.7 are the same as those to the known convergence rates for EI in [16] with no constraints. Meanwhile, the rates in Theorem 3.8 is an improvement over those of [16] and to the best of our knowledge, are the current best convergence rates for CEI (and EI). To achieve this, we applied techniques from recent regret bound analysis on noise-free UCB [22] that allows us to use maximum information gain to bound the sum of $\sigma_{t}^{f}({\bm{x}}_{t+1})$. Then, we use our techniques in the proof of Theorem 3.4 to bound an individual $\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$. In [16], the $\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$ is bounded by the Taylor expansion of the kernel functions. Therefore, the rates of decrease are limited to quadratic terms for both SE and Matérn kernels, since their Taylor expansions around 00 for $\left\lVert{{\bm{x}}-{\bm{x}}^{\prime}}\right\rVert_{2}$ are quadratic at best. On the other hand, maximum information gain has been a focus of research for years, leading to tight bounds on $\gamma_{t}^{f}$ that take advantages of the spectral properties of the kernels [25]. Hence, using $\gamma_{t}^{f}$ to bound $\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$ produces a faster rate. As the open question raised in [26] gets answered, further improvement of the convergence rates is possible, e.g., using techniques from [27].
3.2 Simple regret upper bound under Bayesian assumptions
In this section, we present the simple regret upper bound for CEI under the Bayesian assumptions. We again use the maximum information gain to derive the simple regret upper bound.
@@ 3.2 Simple regret upper bound under Bayesian assumptions
The objective function $f$ is sampled from $GP(0,k_{f}({\bm{x}},{\bm{x}}^{\prime}))$. The constraint function $c$ is sampled from $GP(0,k_{c}({\bm{x}},{\bm{x}}^{\prime}))$. The kernels satisfy $k_{f}({\bm{x}},{\bm{x}}^{\prime})\leq 1$, $k_{f}({\bm{x}},{\bm{x}})=1$, $k_{c}({\bm{x}},{\bm{x}}^{\prime})\leq 1$, and $k_{c}({\bm{x}},{\bm{x}})=1$. Moreover, the bound constraint set $C$ is compact.
Denote the probability operator as $\mathbb{P}$. The objective function $f$ is assumed to be Lipschitz continuous under the Bayesian setting, similar to [28].
Assumption 3.11.
There exists constant $L>0$ such that the following holds
$|f({\bm{x}})-f({\bm{x}}^{\prime})|\leq L\left\lVert{{\bm{x}}-{\bm{x}}^{\prime}}\right\rVert,$ (16)
for $\forall{\bm{x}},{\bm{x}}^{\prime}\in C$.
In the remaining of this section we will work under Assumption 3.10 and 3.11.
@@
In addition to the challenges in the frequentist setting, the bounds associated with the Bayesian setting presents further technical challenges, as they are not available in current literature, to the best of our knowledge. Starting from the bounds on $|f({\bm{x}})-\mu_{t}^{f}({\bm{x}})|$, we derive the bounds on $|I_{t}^{f}({\bm{x}})-EI_{t}^{f}({\bm{x}})|$, an important step towards the bound on $r_{t}$ Noticeably, under the Bayesian setting, the bounds are satisfied with a given probability $1-\delta$, where $\delta\in(0,1)$. Hence, the simple regret upper bounds are presented with a probability as well.
@@
Let $\beta=2\log(6c_{\alpha}/\delta)$, where $c_{\alpha}=\frac{1+2\pi}{2\pi}$. Under Assumption 3.10 and 3.11, the CEI algorithm leads to the simple regret upper bound
$r_{t}\leq$ $c_{\tau}(\beta)\frac{1}{\Phi(-\sqrt{\beta})}\left[\frac{4M_{f}}{t-2}+(0.4+\beta^{1/2})\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\right],$ (17)
for some $t_{k}\in[\frac{t}{2}-1,t]$, $c_{\tau}(\beta)=\frac{\tau(\beta^{1/2})}{\tau(-\beta^{1/2})}$, with probability $\geq 1-\delta$.
The convergence rate is given in the next theorem.
@@
Let $\beta=2\log(6c_{\alpha}/\delta)$, where $c_{\alpha}=\frac{1+2\pi}{2\pi}$. Under Assumption 3.10 and 3.11, the CEI algorithm leads to the convergence rates of
@@
Recall that $f$ and $c$ are assumed conditionally independent in CEI. We start from the bound on the difference between function value and posterior mean for $f$: $|f({\bm{x}})-\mu_{t}^{f}({\bm{x}})|\leq\beta^{1/2}\sigma_{t}^{f}({\bm{x}})$, with probability $\geq 1-\delta$, where $\beta=2\log(1/\delta)$, as in Lemma C.2. Similarly, $|c({\bm{x}})-\mu_{t}^{c}({\bm{x}})|\leq\beta^{1/2}\sigma_{t}^{c}({\bm{x}})$ with probability $\geq 1-\delta$. These are well-known results [28]. Then, we derive the subsequent bounds $|I_{t}^{f}({\bm{x}})-EI_{t}^{f}({\bm{x}})|\leq\sqrt{\beta}\sigma_{t}^{f}({\bm{x}})$, where $\beta=\max\{1.44,2\log(c_{\alpha}/\delta)\}$ and $c_{\alpha}=\frac{1+2\pi}{2\pi}$ with probability $\geq 1-\delta$ (Lemma C.4). Then, we prove the relationship in Lemma C.5 that $I_{t}^{f}({\bm{x}})\leq\frac{\tau(\sqrt{\beta})}{\tau(-\sqrt{\beta})}EI_{t}^{f}({\bm{x}})$ with probability $\geq 1-\delta$. We can now follow the general analysis framework in Section 3.1 and Theorem 3.4 to obtain the simple regret upper bound under Bayesian assumptions.
@@
Comparing Theorem 3.8 to Theorem 3.13, the convergence rates in the frequentist and Bayesian settings are the same. This is because simple regret focuses on the best feasible solution $f_{t}^{+}$ and thus the parameters in Theorem 3.4 and 3.12 do not depend on $t$. That is, $c_{\tau B}$, $B_{c}$, $B_{f}$, $M_{f}$, and $\beta$ do not depend on $t$. However, we note that $\beta$ depends on the desired probability $\delta$.
@@
As mentioned in Section 1, our results can be readily applied to CEI with multiple constraints for both frequentist and Bayesian settings. Consider $m$ constraints $c_{i}({\bm{x}})\leq 0,i=1,\dots,m$. Assuming conditional independence of the constraints, the CEI function is $EI^{C}_{t}({\bm{x}})=\Pi_{i=1}^{m}P_{t}^{i}({\bm{x}})EI_{t}^{f}({\bm{x}})=\Pi_{i=1}^{m}\Phi\left(-\frac{\mu_{t}^{c_{i}}({\bm{x}})}{\sigma_{t}^{c_{i}}({\bm{x}})}\right)EI_{t}^{f}({\bm{x}})$, where $P_{t}^{i}$ is the probability of feasibility function of constraint $c_{i}({\bm{x}})\leq 0$, and $\mu_{t}^{c_{i}}({\bm{x}})$ and $\sigma_{t}^{c_{i}}({\bm{x}})$ are the posterior mean and standard deviation for $c_{i}$, respectively. By making the assumption that each constraint function lies in its corresponding RKHS of the kernel $k_{c_{i}}$, we have $|c_{i}({\bm{x}})-\mu_{t}^{c_{i}}({\bm{x}})|\leq B_{c_{i}}\sigma_{t}^{c_{i}}({\bm{x}})$, where $B_{c_{i}}$ is the upper bound of RKHS norm associated with kernel $k_{c_{i}}$ and function $c_{i}$. We can then apply the analysis framework in this paper to obtain an upper bound similar to that of Theorem 3.4, where the term $\frac{1}{\Phi(-B_{c})}$ is replaced with $\Pi_{i=1}^{m}\frac{1}{\Phi(-B_{c_{i}})}$.
4 Numerical examples
@@
Although this paper is primarily theoretical, we conduct numerical experiments to support the theoretical results. Specifically, we apply the CEI algorithm on two synthetic examples that are randomly generated and five benchmark problems that are commonly used in the CBO literature. These numerical experiments are not intended to demonstrate superior performance over the state-of-the-art CBO algorithms. Instead, they serve as empirical evidence for the theoretical analysis presented in this work.
We first consider an example where both the objective and constraint functions come from the RKHS of their respective kernels. The functions are generated in a similar approach to [29]. The domain of the problem is $[0,1]^{2}$ and both $f$ and $c$ use SE kernels with a length scale of $0.2$. We randomly generate $100$ basis points, and sample from a multi-variate Gaussian distribution at each point. The posterior means of these samples define the objective and constraint functions, whose contours are shown in Figure 2. This example corresponds to the frequentist setting.
Example 2 considers functions sampled from GP with a Matérn kernel (length scale $0.2$ and $\nu=2.5$), following an approach similar to [28]. We uniformly choose $30\times 30$ points in $[0,1]^{2}$, and sample randomly from a multi-variate zero-mean Gaussian distribution. The contours of the objective and constraint functions are shown in Figure 3. This example corresponds to the Bayesian setting.
(a) Example 1: functions in RKHS of SE kernel.
(b) Example 2: functions sampled from GP with Matérn kernel and $\nu=2.5$
(c) Example 3: a small feasible region problem
(d) Example 4: a multiple-constraints problem
(e) Example 5: a four-dimensional problem
(f) Example 6: Hartmann6 function
(g) Example 7: Banana function (The plot is shown in log scale due to the large range of values.)
Figure 1: Simple regret of CEI for seven examples.

### v2

@@ 3.1.1 Improved simple regret upper bound under frequentist assumptions
The proof follows similar steps to that of Theorem 3.3 but further bounds $\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$ using $\gamma_{t}^{f}$. To do so, we first recognize that the bound using $\gamma_{t}^{f}$ (Lemma A.3) is established in the noisy case where the posterior standard deviation has a different form as in (18). Using Lemma A.4, we can establish that the noise-free posterior standard deviation also satisfies $\sum_{i=0}^{t-1}\sigma_{i}^{f}({\bm{x}}_{i+1})\leq\sqrt{C_{\gamma}t\gamma_{t}^{f}}$. Then, from Lemma A.5, we can find a small enough $\sigma_{i}^{f}({\bm{x}}_{i+1})$. Specifically, choose $k=[t/3]$, where $[x]$ denotes the largest integer smaller than $x$. Thus, we have $3k\leq t\leq 3(k+1)$. Then, there exists $k\leq t_{k}\leq 3k$ such that $f^{+}_{t_{k}}-f^{+}_{t_{k}+1}\leq\frac{2B_{f}}{k}$ and $\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\leq\frac{\sqrt{t\gamma_{t}^{f}}}{k}$. The rest of the proof follows from that of Theorem 3.7.
@@
As mentioned above, the rates in Corollary 3.6 are the same as the known convergence rates for EI in Bull (2011). Meanwhile, the rates in Theorem 3.7 is an improvement over those of Bull (2011) for SE kernel with $d\geq 3$ and Matérn kernels with $d\geq 3,\nu\geq\frac{d}{d-2}$. To achieve this, we applied techniques from regret bound analysis on noise-free UCB (Lyu et al., 2019) that allows us to use maximum information gain to bound the sum of $\sigma_{t}^{f}({\bm{x}}_{t+1})$. Then, we use our techniques in the proof of Theorem 3.3 to bound an individual $\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$. In Bull (2011), the $\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$ is bounded by the Taylor expansion of the kernel functions. Therefore, the rates of decrease are limited to quadratic terms for both SE and Matérn kernels, since their Taylor expansions around 0 for $\left\lVert{{\bm{x}}-{\bm{x}}^{\prime}}\right\rVert_{2}$ are quadratic at best. On the other hand, maximum information gain can lead to tight bounds on $\gamma_{t}^{f}$ that take advantages of the spectral properties of the kernels (Vakili et al., 2021; Iwazaki, 2025b). Hence, using $\gamma_{t}^{f}$ to bound $\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$ can produce a faster rate. As the open question raised in Vakili (2022) gets answered, further improvement of the convergence rates is possible, e.g., using techniques from Iwazaki (2025a).
3.2 Simple regret upper bound under Bayesian objective assumption
In this section, we present the simple regret upper bound for CEI under the Bayesian objective assumptions. We again use the maximum information gain to derive the simple regret upper bound.
@@ 3.2 Simple regret upper bound under Bayesian objective assumption
The bound constraint set $C\subset[0,r]^{d}$ is compact and convex. The objective function $f$ is sampled from $GP(0,k_{f}({\bm{x}},{\bm{x}}^{\prime}))$. Further, the objective function $f$ is assumed to be Lipschitz continuous (of 1-norm) with Lipschitz constant $L_{f}$ with probability $\geq 1-da_{f}e^{L_{f}^{2}/b_{f}^{2}}$ for some constants $a_{f}>0$ and $b_{f}>0$. The kernels satisfy $k_{f}({\bm{x}},{\bm{x}}^{\prime})\leq 1$ and $k_{f}({\bm{x}},{\bm{x}})=1$. The constraint function $c$ remains in the RKHS of $k_{c}$, similarly to the frequentist setting.
In the remaining of this section we will work under Assumption 3.9.
@@
In addition to the challenges in the frequentist setting, the bounds on EI in the Bayesian setting are not available in current literature, to the best of our knowledge. Starting from the confidence interval on $|f({\bm{x}})-\mu_{t}^{f}({\bm{x}})|$, we derive the bounds on $|I_{t}^{f}({\bm{x}})-EI_{t}^{f}({\bm{x}})|$ with high probability, an important step towards the bound on $r_{t}$ . Noticeably, under the Bayesian setting, the bounds are satisfied with a given probability, e.g., $1-\delta$, where $\delta\in(0,1)$.
@@
Let $\beta=2\log(6c_{\alpha}/\delta)$ and $\beta_{t}=2\log(3\pi_{t}/\delta)$, where $c_{\alpha}=\frac{1+2\pi}{2\pi}$ and $\pi_{t}=\frac{\pi^{2}t^{2}}{6}$. Under Assumption 3.9, the CEI algorithm leads to the simple regret upper bound
$r_{t}\leq$ $c_{\tau}(\beta)\frac{1}{\Phi(-B_{c})}\left[\frac{4M_{f}}{t-2}+\frac{2\beta^{1/2}_{t}}{t-2}\sqrt{C_{\gamma}t\gamma_{t}^{f}}+(0.4+\beta^{1/2})\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\right],$ (16)
for some $t_{k}\in[\frac{t}{2}-1,t]$, $c_{\tau}(\beta)=\frac{\tau(\beta^{1/2})}{\tau(-\beta^{1/2})}$, and constant $M_{f}>0$ with probability $\geq 1-\delta$.
The constant $M_{f}$ is from Lemma C.1. The convergence rate is given in the next theorem.
@@
Let $\beta=2\log(6c_{\alpha}/\delta)$ and $\beta_{t}=2\log(3\pi_{t}/\delta)$, where $c_{\alpha}=\frac{1+2\pi}{2\pi}$ and $\pi_{t}=\frac{\pi^{2}t^{2}}{6}$. Under Assumption 3.9, the CEI algorithm leads to the convergence rates of
$\mathcal{O}\left(t^{-\frac{1}{2}}\log^{\frac{d+2}{2}}(t)\right)\ \text{and }\ \mathcal{O}\left(t^{\frac{-\nu}{2\nu+d}}\log^{\frac{2\nu+0.5d}{2\nu+d}}(t)\right),$ (17)
@@
Recall that $f$ and $c$ are assumed conditionally independent in CEI. We start from the bound on the confidence interval for $f$: $|f({\bm{x}})-\mu_{t}^{f}({\bm{x}})|\leq\beta^{1/2}\sigma_{t}^{f}({\bm{x}})$, with probability $\geq 1-\delta$, where $\beta=2\log(1/\delta)$, as in Lemma C.2. The confidence interval of $c$ remains the same as in the frequentist setting. These are well-known results (Srinivas et al., 2009). Then, we derive the subsequent bounds $|I_{t}^{f}({\bm{x}})-EI_{t}^{f}({\bm{x}})|\leq\sqrt{\beta}\sigma_{t}^{f}({\bm{x}})$, where $\beta=\max\{1.44,2\log(c_{\alpha}/\delta)\}$ and $c_{\alpha}=\frac{1+2\pi}{2\pi}$ with probability $\geq 1-\delta$ (Lemma C.5). Then, we prove the relationship in Lemma C.6 that $I_{t}^{f}({\bm{x}})\leq\frac{\tau(\sqrt{\beta})}{\tau(-\sqrt{\beta})}EI_{t}^{f}({\bm{x}})$ with probability $\geq 1-\delta$. We can now follow the general analysis framework in Section 3.1 and Theorem 3.3 to obtain the simple regret upper bound under Bayesian objective assumptions, while choosing $t_{k}$ with a more defined criterion.
@@
Comparing Theorem 3.7 to Theorem 3.11, the convergence rates in the frequentist and Bayesian settings are the same except for a $\log^{1/2}(t)$ term. This is partially because simple regret focuses on the best feasible solution $f_{t}^{+}$ and thus many of the parameters in Theorem 3.3 and 3.10 do not depend on $t$.
@@
As mentioned in Section 1, our results can be readily applied to CEI with multiple constraints for both frequentist and Bayesian settings. Consider $m$ constraints $c_{i}({\bm{x}})\leq 0,i=1,\dots,m$. Assuming conditional independence of the constraints, the CEI function is $EI^{C}_{t}({\bm{x}})=\Pi_{i=1}^{m}P_{t}^{i}({\bm{x}})EI_{t}^{f}({\bm{x}})=\Pi_{i=1}^{m}\Phi\left(\frac{-\mu_{t}^{c_{i}}({\bm{x}})}{\sigma_{t}^{c_{i}}({\bm{x}})}\right)EI_{t}^{f}({\bm{x}})$, where $P_{t}^{i}$ is the probability of feasibility function of constraint $c_{i}({\bm{x}})\leq 0$, and $\mu_{t}^{c_{i}}({\bm{x}})$ and $\sigma_{t}^{c_{i}}({\bm{x}})$ are the posterior mean and standard deviation for $c_{i}$, respectively. By making the assumption that each constraint function lies in its corresponding RKHS of the kernel $k_{c_{i}}$, we have $|c_{i}({\bm{x}})-\mu_{t}^{c_{i}}({\bm{x}})|\leq B_{c_{i}}\sigma_{t}^{c_{i}}({\bm{x}})$, where $B_{c_{i}}$ is the upper bound of RKHS norm associated with kernel $k_{c_{i}}$ and function $c_{i}$. We can then apply the analysis framework in this paper to obtain an upper bound similar to that of Theorem 3.3, where the term $\frac{1}{\Phi(-B_{c})}$ is replaced with $\Pi_{i=1}^{m}\frac{1}{\Phi(-B_{c_{i}})}$. We note that in the Bayesian objective setting, to ensure probability $1-\delta$, the parameter $\beta$ needs to increase with the number of constraints as well, e.g., $\beta=2\log((m+5)c_{\alpha}/\delta)$.
Remark 3.14 (Extension to the noisy setting).
Extending our analysis to the noisy setting is non-trivial, and we discuss the associated challenges for noisy objective and constraint functions separately. A noisy constraint function introduces additional complications in defining feasibility. If only noisy observations of the constraint values are available, the notion of a feasible sample and the definition of $f_{t}^{+}$ becomes ambiguous. As a result, major modifications to the CEI algorithm are required to appropriately handle the uncertainty introduced by noise.
For the noisy objective function, CEI can be adapted similarly to the noisy EI formulation by treating the best feasible noisy observation as the incumbent. However, to the best of our knowledge, a theoretical guarantee on the simple regret bound for the noisy unconstrained setting remains unavailable. Recent work by Wang et al. (2025) provides a framework for deriving noisy simple regret bounds based on the best observed value, $r_{t}^{s}=y_{t}^{+}-f({\bm{x}}_{t})$, which can be extended to CEI. Specifically, by defining $r_{t}^{s}$ as the simple regret for CEI with $y_{t}^{+}$ denoting the best feasible noisy observation, a similar proof strategy as in Theorem 3.10 yields an analogous upper bound. In the Bayesian setting with i.i.d. Gaussian noise on the objective and noise-free constraint observations, the convergence rate of the upper bound on $r_{t}^{s}$ can be obtained. However, we note that given the noise, $r_{t}^{s}$ is possibly negative.
Remark 3.15 (Infeasible initial sample).
It is well known that CEI requires initial feasible sample (Gardner et al., 2014). That is, $f_{t}^{+}$ exists from the initial samples so that the CEI calculation can proceed. Methods proposed to address this issue typically employ separate strategies when no feasible samples are available and revert to the standard CEI formulation once feasibility is established (Lin et al., 2024; Letham et al., 2019). In addition, introducing a tolerance parameter in the constraint can further mitigate this problem by allowing near-feasible points when the degree of violation is small.
Remark 3.16 (Tolerance in constraints).
In gradient-based optimization methods, a tolerance for constraint violation is often used to improve the performance and flexibility of algorithms Wächter and Biegler (2006); Nocedal and Wright (2006). Motivated by this, we introduce a tolerance parameter $\lambda\geq 0$, where a point $\bm{x}$} is considered feasible if $c({\bm{x}})\leq\lambda$ and infeasible otherwise. The corresponding CEI with tolerance is defined as $EI^{C}_{t}({\bm{x}},\lambda)=P_{t}({\bm{x}},\lambda)EI_{t}^{f}({\bm{x}})=\Phi\left(\frac{\lambda-\mu_{t}^{c}({\bm{x}})}{\sigma_{t}^{c}({\bm{x}})}\right)EI_{t}^{f}({\bm{x}}).$ Clearly, the standard CEI formulation is recovered when $\lambda=0$.
The simple regret bound is affected by $\lambda$ and should lead to $\frac{1}{\Phi(\lambda-B_{c})}$ in place of $\frac{1}{\Phi(-B_{c})}$. In fact, we can replace $\frac{1}{\Phi(\lambda-B_{c})}$ with $1/{\Phi\left(\frac{\lambda}{\sigma^{c}_{t_{k}}(x^{*})}-B_{c}\right)}$, which is time-varying. One can follow the proof of Theorem 3.3 to obtain this term, which emerges from the confidence interval of $c$ at $t_{k}$, $x^{*}$ and $c(x^{*})\leq 0$.
As the sample iteration increases, the inclusion of $\sigma^{c}_{t_{k}}(x^{*})$ is important in balancing $B_{c}$ that can lead to a large simple regret upper bound. We explain the intuition below. As $t\to\infty$, $t_{k}\to\infty$ and $k\to\infty$. We know $\sigma_{t}^{f}({\bm{x}}_{t+1})\to 0$, and hence $\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\to 0$ and $r_{t}\to 0$. That is, the simple regret upper bound of CEI with $\lambda>0$ converges. Thus, ${\bm{x}}_{t}$ approaches at least one of the optimal solutions. Suppose without losing generality, ${\bm{x}}_{t}\to{\bm{x}}^{*}$. Then, by definition $\sigma_{t_{k}}^{c}({\bm{x}}^{*})\to 0$. Consequently, we should have $\frac{\lambda}{\sigma^{c}_{t_{k}}(x^{*})}\rightarrow\infty$ for $\lambda>0$. Then, we have $\Phi\left(\frac{\lambda}{\sigma^{c}_{t_{k}}(x^{*})}\right)\to 1$. Therefore, $1/{\Phi(\frac{\lambda}{\sigma^{c}_{t_{k}}(x^{*})}-B_{c})}\to 1$ and $B_{c}$ does not affect the simple regret upper bound asymptotically. We note that the convergence rate of CEI with tolerance remains similar since it is dominated by the maximum information gain of $f$.
4 Numerical experiments
@@
Although this paper is primarily theoretical, we conduct numerical experiments to support the theoretical results. We apply the CEI algorithm to eight synthetic problems that are randomly generated from RKHS of kernels and GP priors, and five benchmark problems commonly used in the CBO literature. These numerical experiments are not intended to demonstrate superior performance over the state-of-the-art CBO algorithms. Instead, they serve as empirical evidence for the theoretical analysis presented in this work. All experiments are conducted on M1 (16GB memory) [^1] Codes are available in https://github.com/Haowei-Wang/Convergence-Rates-of-Constrained-Expected-Improvement..
4.1 Synthetic problems
In this section, we study objective and constraint functions drawn from reproducing kernel Hilbert spaces (RKHSs) as well as from Gaussian process priors with Matérn ($\nu=2.5$) and squared exponential (SE) kernels, across input dimensions $d\in{2,4}$. The domain is the hypercube $0,1]^{d}$. For RKHS cases (the frequentist setting), the functions are generated with a similar approach to Chowdhury and Gopalan (2017). Specifically, both objective $f(x)$ and constraint functions $c(x)$ are generated by sampling from the RKHS associated with a chosen kernel (Matérn/SE kernels with a length scale of $0.2$). Each function is constructed as a weighted sum of kernel evaluations at $100$ randomly selected basis points, with weights drawn from a standard normal distribution. Formally, the function takes the form $f(x)=\sum_{i=1}^{n}\alpha_{i}k(x,X_{i})$ , where $k$ is the kernel, $X_{i}$ are basis points, and $\alpha_{i}$ are random coefficients; $c(x)$ is generated similarly. For the GP cases (the Bayesian setting), the functions are generated with an approach similar to Srinivas et al. (2009). Specifically, we uniformly choose 1000 points in the design space and sample randomly from a multivariate Gaussian distribution defined by the GP prior with the chosen kernel.
For each synthetic problem, we conducted 100 independent trials. The number of initial design is set to $10d$, and 50 optimization iterations were performed for all cases. We plotted the log-log curve of simple regret against the number of iterations in Figure 1. In all cases, we consistently observed sublinear convergence patterns, which align well with our theoretical guarantees.
(a) RKHS, Matérn, $d=2$
(b) RKHS, Matérn, $d=4$
(c) RKHS, SE, $d=2$
(d) RKHS, SE, $d=4$
(e) GP, Matérn, $d=2$
(f) GP, Matérn, $d=4$
(g) GP, SE, $d=2$
(h) GP, SE, $d=4$
Figure 1: The log-log plots for simple regret vs optimization iterations of CEI for the synthetic problems.
4.2 Test problems